Session: Jun 2007 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Fri Jun 1 14:32:53 2007) Success :
No problem during observations

Shanghai (Bo Xia at Fri Jun 1 14:34:25 2007) Success :

Urumqi (Zhang hua at Fri Jun 1 15:02:17 2007) Minor failures :
The compressor of C-band doesn't working.we used room temperature receiver. Otherwise,It is rain. No known problems.

Onsala ( Michael Lindqvist at Fri Jun 1 16:26:08 2007) Success :
Slightly higher SEFD in LCP than usual (normally the SEFD in LCP and RPC are about the same). Otherwise no known problems.

Noto (P. Cassaro at Sun Jun 3 09:58:37 2007) Success :

Torun (K. Borkowski at Mon Jun 4 07:11:02 2007) Success :
No known problems.

Effelsberg (A. Polatidis at Mon Jun 4 07:59:16 2007) Minor failures :
VC05 still not functioning. Seemed to be OK, but no fringes in the FTP test. We are investigating what could have been wrong.

Westerbork (Tony Foley at Mon Jun 4 08:35:55 2007) Minor failures :
Wb used in phased array mode. Data check at 12:33 gave some errors. Others OK